# Adversarial Attacks on Data Attribution

### v1

@@ 6 Experiments
In this section, we empirically evaluate the effectiveness of the two proposed methods, Shadow Attack and Outlier Attack.
@@ 6.1 Experimental Setup
We consider two machine learning tasks, image classification, and text generation. For image classification, we experiment on MNIST (LeCun, 1998) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ two data attribution methods, Influence Function (Koh & Liang, 2017) and TRAK (Park et al., 2023). For text generation, we conduct experiments on NanoGPT trained on the Shakespeare dataset (Karpathy, 2022), with TRAK as the data attribution method. For the image classification settings, we evaluate both Shadow Attack and Outlier Attack on image classification tasks, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. The experimental settings are summarized in Table 1.
Data contribution workflow.
For the image classification settings, we set $|Z_{0}|=10000,|Z_{1}^{a}|=100$, and $|Z_{1}|=11000$. This setup simulates the following data contribution workflow: At $t=0$ when there is no Adversary, and $10000$ training points are used to train the model $\mathcal{T}(Z_{0})$. At $t=1$, the Adversary contributes $100$ perturbed training points and other Data Providers contribute $900$ new training data points. Together with the previous $10000$ training points, a total of $11000$ training points are used to train the model $\mathcal{T}(Z_{1})$. The text generation setting follows a similar workflow with $|Z_{0}|=4706,|Z_{1}^{a}|=20$, and $|Z_{1}|=6274$.
Evaluation metrics.
We evaluate the proposed attack methods using two metrics. The first is the Compensation Share, as defined in Eq. 1. For each attack method, we calculate the Compensation Share respectively when the original dataset without perturbation is contributed as $Z_{1}^{a}$ (Original) and when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$ (Manipulated). The increase of the Compensation Share after the perturbation (as measured by the Ratio) reflects the effectiveness of the attack.
To gain a more detailed understanding of how the adversarial perturbations affect the top-$k$ influential data points for individual validation data points in $V_{1}$ we introduce the second evaluation metric named Fraction of Change. For each validation data point $v\in V_{1}$, we measure how many data points in $Z_{1}^{a}$ appear in the top-$k$ influential data points for $v$. In comparison to when the original dataset without perturbation is contributed as $Z_{1}^{a}$, we calculate the fraction of validation data points in $V_{1}$ that contain more data points from $Z_{1}^{a}$ in the top-$k$ influential data points when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$. Similarly, we calculate the fraction of validation points that contain the same number of or fewer data points from $Z_{1}^{a}$ after perturbation. We report the three fractions under the categories More, Tied, and Fewer for the Fraction of Change metric. The higher fraction for the More category indicates that the attack method influences the validation data points more broadly.
For both metrics, we set $k=100$ when counting the top-$k$ influential data points.
Table 1: Summary of experimental settings with various tasks, datasets, target models, and data attribution methods.
@@
(b) Image Classification MNIST CNN TRAK
(c) Image Classification CIFAR-10 ResNet-18 TRAK
(d) Text Generation Shakespeare NanoGPT TRAK
@@ 6.2 Experimental Results: Shadow Attack
The results of the proposed Shadow Attack method are shown in Table 2. We conduct experiments on all three image classification settings outlined in Table 1. For each of the settings, we first consider the case that the shadow models have the same architecture as the target model, as shown in the first three rows of Table 2. Additionally, for the third setting where the target model is ResNet-18, we further consider using ResNet-9 as the shadow model, which simulates scenarios where the Adversary’s knowledge about the training algorithm is limited.
Table 2: Results of the Shadow Attack method. The target models used for evaluation in each setting are listed in Table 1 while the shadow models used in the attack are listed under the column “Shadow Model”. The proportion of the data contributed by the Adversary relative to the full dataset at $t=1$ is shown under the column “$|Z_{1}^{a}|/|Z_{1}|$”. The Original dataset consists of random samples without perturbation, so its Compensation Share aligns closely with $|Z_{1}^{a}|/|Z_{1}|$. Overall, a higher Ratio indicates a more effective attack. A higher fraction under More means that the attack influences the validation data points more broadly.
@@
(a) LR 0.0098 0.0098 0.0477 456.1% 0.955 0.038 0.007
(b) CNN 0.0098 0.0112 0.0467 417.0% 0.781 0.195 0.024
(c) ResNet-18 0.0098 0.0095 0.0213 217.3% 0.655 0.259 0.086
(c) ResNet-9 0.0098 0.0095 0.0196 206.3% 0.622 0.310 0.068
The Shadow Attack method is highly effective in increasing the Compensation Share across all the settings. The Ratio of the Manipulated to the Original ranges from $206.3\%$ to $456.1\%,$ representing a substantial increase. Notably, the last row corresponds to the setup where the shadow models have the ResNet-9 architecture while the target model is a ResNet-18, and the Shadow Attack remains significantly effective ($206.3\%,$) although being slightly worse than the case where shadow model and target model shares the same architecture (the third row, $217.3\%.$)
@@ 6.3 Experimental Results: Outlier Attack
On experimental settings (a) and (b), the Outlier Attack performs even better than the Shadow Attack in terms of both Compensation Share and Fraction of Change, achieving a higher Ratio and a larger fraction under More. Across all four settings, the Ratio ranges from $185.2\%$ to $643.9\%,$ demonstrating the exceptional effectiveness of the Outlier Attack. The results from setting (d), the text generation setting, further highlight the generalizability of the proposed method.
Table 3: Results of the Outlier Attack method. The black-box adversarial attack methods for generating adversarial examples are listed under the column “Attack Method”. See Table 2 for more context.
@@
(a) ZOO 0.0098 0.0098 0.0631 643.9% 0.980 0.017 0.003
(b) Simba 0.0098 0.0112 0.0668 596.4% 0.799 0.173 0.028
(c) Simba 0.0098 0.0095 0.0176 185.2% 0.562 0.354 0.084
(d) TextFooler 0.0031 0.0035 0.0092 262.9% 0.392 0.461 0.147
@@ 6.4 Baseline Reference
To better understand the significance of the results in Table 2 and Table 3, we present experimental results for a baseline random perturbation method in Table 4. We conduct experiments on the three settings of image classification, where the baseline method applies pixel-wise random perturbation to the images, with the perturbation budget matched to that of the proposed attack methods. As shown in the results, the Compensation Share of the dataset with Random Perturbation is nearly identical to that of the Original. The fraction under More is also close to that under Fewer. These results further emphasize that the effectiveness of the proposed attack methods comes from the careful design of adversarial perturbation.
@@ 6.5 Visualization of the Perturbation
Finally, we visualize two examples of the adversarially perturbed images in Figure 3 and Figure 4, respectively from MNIST and CIFAR-10. As can be seen, the perturbations are barely visible to human eyes. However, these perturbations make the images significantly more influential for a large number of validation data points.
(a) Original image. Influential for 00 validation data points.
(b) Image after Shadow Attack. Influential for $75$ validation data points.
(c) Image after Outlier Attack. Influential for $105$ validation data points.
Figure 3: Visualization of an image from MNIST, before and after manipulation.
(a) Original image. Influential for $1$ validation data points.
(b) Image after Shadow Attack. Influential for $38$ validation data points.
(c) Image after Outlier Attack. Influential for $29$ validation data points.
Figure 4: Visualization of an image from CIFAR-10, before and after manipulation.

### v2

@@ 6.1 Experimental setup
We consider two machine-learning tasks: image classification and text generation. For image classification, we experiment on MNIST (LeCun, 1998), Digits (Jiang et al., 2023) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Multi-layer Perceptron (MLP), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ three popular data attribution methods, including Influence Function (Koh & Liang, 2017), TRAK (Park et al., 2023), and Data Shapley (Ghorbani & Zou, 2019). For text generation, we conduct experiments on NanoGPT (Karpathy, 2022) trained on the Shakespeare dataset (Karpathy, 2015), with TRAK as the data attribution method. Finally, for the image classification settings, we evaluate both Shadow Attack and Outlier Attack, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. The experimental settings are summarized in Table 1.
Table 1: Summary of the experimental settings.
@@
(b) Image Classification Digits MLP Data Shapley
(c) Image Classification MNIST CNN TRAK
(d) Image Classification CIFAR-10 ResNet-18 TRAK
(e) Text Generation Shakespeare NanoGPT TRAK
Data Contribution Workflow.
For the image classification settings (a), (c), and (d), we set $|Z_{0}|=10000,|Z_{1}^{a}|=100$, and $|Z_{1}|=11000$. This setup simulates the following data contribution workflow: At $t=0$ when there is no Adversary, and $10000$ training points are used to train the model $\mathcal{T}(Z_{0})$. At $t=1$, the Adversary contributes $100$ perturbed training points and other Data Providers contribute $900$ new training data points. Together with the previous $10000$ training points, a total of $11000$ training points are used to train the model $\mathcal{T}(Z_{1})$. For image classification setting (b), due to the size of the dataset, we set $|Z_{0}|=1100$, $|Z_{1}^{a}|=30$ and $|Z_{1}|=1100$ for Outlier Attack, $|Z_{0}|=800$, $|Z_{1}|^{a}=30$ and $|Z_{1}|=850$ for Shadow Attack. The text generation setting follows a similar workflow with $|Z_{0}|=4706$, $|Z_{1}^{a}|=20$, and $|Z_{1}|=6274$.
Evaluation Metrics.
We evaluate the proposed attack methods using two metrics, where we set $k=100$ when counting the top-$k$ influential data points in both cases. The first is the Compensation Share (Eq. 1) where for each attack method, we calculate $c(Z_{1}^{a})$ respectively when the original dataset without perturbation is contributed as $Z_{1}^{a}$ (Original) and when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$ (Manipulated). The increase of $c(Z_{1}^{a})$ after the perturbation measured by the Ratio reflects the effectiveness of the attack.
To gain a more refined understanding of how the adversarial perturbations affect the top-$k$ influential data points for individual validation data points in $V_{1}$, we consider the second evaluation metric named Fraction of Change. For each validation data point $v\in V_{1}$, it measures how many data points in $Z_{1}^{a}$ appear in the top-$k$ influential data points for $v$. In comparison to when the original dataset without perturbation is contributed as $Z_{1}^{a}$, we calculate the fraction of validation data points in $V_{1}$ that contain more data points from $Z_{1}^{a}$ in the top-$k$ influential data points when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$. Similarly, we calculate the fraction of validation points that contain the same number of or fewer data points from $Z_{1}^{a}$ after perturbation. We report the three fractions under the categories More, Tied, and Fewer, where the higher fraction for the More category indicates that the attack method influences the validation data points more broadly.
@@ 6.2 Experimental results: Shadow Attack
The results of the proposed Shadow Attack method are shown in Table 2. We conduct experiments on all three image classification settings outlined in Table 1. For each of the settings, we first consider the case that the shadow models have the same architecture as the target model, as shown in the first three rows of Table 2. Additionally, for setting (d) where the target model is ResNet-18 (forth row), we further consider using ResNet-9 as the shadow model (last row), which simulates scenarios where the Adversary’s knowledge about the training algorithm is limited.
Table 2: Results of the Shadow Attack method. The target models used for evaluation in each setting are listed in Table 1 while the shadow models used in the attack are listed under Shadow Model. The proportion $|Z_{1}^{a}|/|Z_{1}|$ of the data contributed by the Adversary relative to the full dataset at $t=1$ is also reported. A higher Ratio indicates a more effective attack, and a higher fraction under More means that the attack influences the validation data points more broadly.
@@
(a) LR $0.0098$ $0.0098$ $0.0477$ $456.1\%$ $0.955$ $0.038$ $0.007$
(b) MLP $0.0352$ $0.0152$ $0.0435$ $286.2\%$ $0.533$ $0.333$ $0.134$
(c) CNN $0.0098$ $0.0112$ $0.0467$ $417.0\%$ $0.781$ $0.195$ $0.024$
(d) ResNet-18 $0.0098$ $0.0095$ $0.0213$ $217.3\%$ $0.655$ $0.259$ $0.086$
(d) ResNet-9 $0.0098$ $0.0095$ $0.0196$ $206.3\%$ $0.622$ $0.310$ $0.068$
The Shadow Attack method is highly effective in increasing the Compensation Share across all the settings. The Ratio of the Manipulated to the Original ranges from $206.3\%$ to $456.1\%,$ representing a substantial increase. Notably, the last row corresponds to the setup where the shadow models have the ResNet-9 architecture while the target model is a ResNet-18, and the Shadow Attack remains significantly effective ($206.3\%,$) although being slightly worse than the case where shadow model and target model shares the same architecture (forth row, $217.3\%.$).
@@ 6.3 Experimental results: Outlier Attack
Table 3: Results of the Outlier Attack method. The black-box adversarial attack methods for generating adversarial examples are listed under Attack Method. See Table 2 for more context.
@@
(a) ZOO $0.0098$ $0.0098$ $0.0631$ $643.9\%$ $0.980$ $0.017$ $0.003$
(b) Simba $0.0250$ $0.0112$ $0.0218$ $194.6\%$ $0.440$ $0.380$ $0.180$
(c) Simba $0.0098$ $0.0112$ $0.0668$ $596.4\%$ $0.799$ $0.173$ $0.028$
(d) Simba $0.0098$ $0.0095$ $0.0176$ $185.2\%$ $0.562$ $0.354$ $0.084$
(e) TextFooler $0.0031$ $0.0035$ $0.0092$ $262.9\%$ $0.392$ $0.461$ $0.147$
On experimental settings (a) and (c), the Outlier Attack performs even better than the Shadow Attack in terms of both Compensation Share and Fraction of Change, achieving a higher Ratio and a larger fraction under More. Across all four settings, the Ratio ranges from $185.2\%$ to $643.9\%,$ demonstrating the exceptional effectiveness of the Outlier Attack. Finally, the results of text generation setting (e) further highlight the applicability of the proposed method to generative AI models.
@@ 6.4 Baseline reference
To better understand the significance of the results in Table 2 and Table 3, we compare them with a baseline random perturbation method in Table 4 on the three image classification settings such that the baseline method applies pixel-wise random perturbation to the images, with the perturbation budget matched to that of the proposed attack methods. As shown in the results, the Compensation Share of the dataset with Random Perturbation is nearly identical to that of the Original. The fraction under More is also close to that under Fewer. These results highlight the effectiveness of the proposed attack methods comes from the careful design of adversarial perturbation.
@@ 6.5 Visualization of the perturbation
Finally, Figure 3 provides visualizations of two examples of the adversarial perturbed images CIFAR-10. In both cases, the perturbations are barely visible to human eyes, however, they are highly effective in terms of the success of the attack.
(a) Original.
Influential for 00 validation data points.
(b) Shadow Attack.
Influential for $75$ validation data points.
(c) Outlier Attack.
Influential for $105$ validation data points.
(d) Original.
Influential for $1$ validation data points.
(e) Shadow Attack.
Influential for $38$ validation data points.
(f) Outlier Attack.
Influential for $29$ validation data points.
Figure 3: Visualization of MNIST (Top) and CIFAR-10 (Bottom), before and after attacks.